OpenStudy (anonymous):

find the vertex: y=5x²-3x+2

OpenStudy (anonymous):

first coordinate of vertex is \[-\frac{b}{2a}=-\frac{-3}{2\times 5}=-\frac{3}{10}\] to find the second coordinate replace x by \[-\frac{3}{10}\] and see what you get

OpenStudy (anonymous):

oopos

Directrix (directrix):

(3/10, 31/20) See graph at http://www.wolframalpha.com/input/?i=y+%3D+5x%C2%B2+-+3x+%2B+2

OpenStudy (anonymous):

\[\frac{3}{10}\] i was off by a minus sign, sorry